OpenStudy (anonymous):

will give medal and become fan! 2nd to last one(:

OpenStudy (anonymous):

@Chelsea04

OpenStudy (anonymous):

@ryan123345 @azolotor

OpenStudy (anonymous):

the eqn is: (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

r is your radius

OpenStudy (jtvatsim):

just take the square root of 64, that means the radius is 8. In general, a circle is written as: (x - A)^2 + (y - B)^2 = r^2 where (A,B) is the center point and r is the radius.

OpenStudy (jtvatsim):

sorry you were already there... my bad Chelsea04 :)

OpenStudy (jtvatsim):

h, k is more standard, but it's late for me, so I'm just making up letters since it doesn't really matter. :)

OpenStudy (jtvatsim):

good luck! later

OpenStudy (anonymous):

so do you do 64 / 8?

OpenStudy (anonymous):

the equation says r^2 but r is your radius so what you need to do is to sqrt 68

OpenStudy (anonymous):

*64

OpenStudy (anonymous):

Oh! I got 8 (: @Chelsea04

OpenStudy (anonymous):

yea, that's your radius :)

OpenStudy (anonymous):

just remember that the equation gives r^2 not r

OpenStudy (anonymous):

okayy! Last questin!

OpenStudy (anonymous):

k!